# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 2 Assumptions and main results
- Parabolic nature: In the aggregated region, $S,$ the systems diverge significantly. While (2.13) exhibits a fully parabolic regime with respect to all three variables $\rho_{1},$ $\rho_{2}$ and $\rho_{1}+\rho_{2},$ equation (2.18) is only parabolic with respect $\rho_{1}+\rho_{2}.$ Indeed, a formal analysis on the behavior of the solution of (2.18) into the three driver regions (assuming again they are sufficiently regular) implies that :
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=f_{2}=0\\ \\ \partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\>\nabla\alpha^{\prime}(\rho_{1})+\rho_{1}V_{1})=f_{1}\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=f_{1}=0,\>\rho_{2}>0\\ \\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\nabla\alpha^{\prime}(\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
$\left\{\begin{array}[]{lll}\partial_{t}\rho_{1}-\nabla\cdot(\sigma_{1}\>\rho_{1}\nabla\alpha^{\prime}(\rho_{1}+\rho_{2})+\rho_{1}V_{1})=f_{1}\\ \partial_{t}\rho_{2}-\nabla\cdot(\sigma_{2}\>\rho_{2}\nabla\alpha^{\prime}(\rho_{1}+\rho_{2})+\rho_{2}V_{2})=f_{2}\end{array}\right.$
@@
Without any external factors influencing the system ($f\equiv 0$ and $V\equiv 0$), in the case where $\sigma_{k}=\alpha_{k}=1,$ for $k=1,2,$ we expect the solution of (2.13) to follow this rule
$\left\{\begin{array}[]{lll}\rho_{1}>0,\>\rho_{2}=0\\ \\ \partial_{t}\rho_{1}-\Delta\beta^{\prime}(\rho_{1})=0\\ -\Delta\eta_{2}=0,\>\eta_{2}\leq\beta^{\prime}(\rho_{1})\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}=0,\>\rho_{2}>0\\ \\ -\Delta\nabla\eta_{1}=0,\>\eta_{1}\leq\beta^{\prime}(\rho_{2})\\ \partial_{t}\rho_{2}-\Delta\beta^{\prime}(\rho_{2})=0\end{array}\right.$
$\left\{\begin{array}[]{lll}\rho_{1}-\rho_{2}=\rho_{01}-\rho_{02}\\ \\ \partial_{t}\rho_{1}-\Delta\beta^{\prime}(\rho_{1}+\rho_{2})=0\\ \partial_{t}\rho_{2}-\Delta\beta^{\prime}(\rho_{1}+\rho_{2})=0\end{array}\right.$
Particularly for homogeneous boundary conditions, we expect $\eta_{1}=0$ (respectively, $\eta_{2}=0$) in region $S_{2}$ (respectively, $S_{1}$). Additionally, given that $\rho_{1}-\rho_{2}=\rho_{01}-\rho_{02}$ in region $S,$ we anticipate $\rho_{1}=\rho_{2}$ outside the combined support of the initial densities $\rho_{01}$ and $\rho_{02}.$
@@
(2.19) $\|f\|_{H^{-1}_{\Gamma_{D_{k}}}}=\min_{\omega\in L^{2}(\Omega)^{N}}\left\{\left(\int_{\Omega}\sigma_{k}\>|\omega|^{2}\right)^{1/2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
@@
$\langle f,g\rangle_{k,\Omega}:=\int_{\Omega}\sigma_{k}\>\phi^{f}\cdot\phi^{g},\quad\hbox{ for any }f,g\>\in H^{-1}_{\Gamma_{D_{k}}}(\Omega)$
where $\phi^{f}$ and $\phi^{g}$ are given by
$\phi^{f}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}$
@@
Now let us consider the space $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),$ endowed with the norm
@@
This is in turn a Hilbert space, and the the associate inner product is given by
@@
(2.20) $\mathcal{E}(\rho):=\left\{\begin{array}[]{ll}\int_{\Omega}\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})&\hbox{ if }\rho_{1},\>\rho_{2},\>\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})\in L^{1}_{+}(\Omega)\\ \\ +\infty&\hbox{ otherwise }.\end{array}\right.$
Thanks to assumptions $(H1)$ and $(H2)$, one sees that $\mathcal{E}$ is convex and l.s.c. in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega).$ This result implies that the sub-differential of $,\displaystyle\mathcal{E},$$\partial\mathcal{E},$ is well-defined and forms a maximal monotone graph within $\left(H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),\|-\|_{H^{-1}_{\Gamma_{D}}}\right).$ Thus we can consider the $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow problem
(2.21) $\left\{\begin{array}[]{ll}\frac{d}{dt}\rho(t)+\partial\mathcal{E}(\rho(t))\ni f(t)&\hbox{ a.e. }t\in(0,T)\\ \\ \rho(0)=\rho_{0}.\end{array}\right.$
Our aim now, is to establish the connection the problems (2.13) and (2.21). This is summarized in the following theorem.
@@
Under the assumptions $g\equiv 0$, $\pi\equiv 0$, $V\equiv 0,$ $f\in BV\left(0,T;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\right)$, let us consider $(\rho,\eta)$ be a weak solution of the system of PDE (2.13). Then, $\rho$ coincides with the unique strong solution of (2.21).
@@
    Theorem 2 bridges a connection between the cross-diffusion system (3.37) and the well-established theory of gradient flows in Hilbert spaces. Furthermore, the specific form of the internal energy (2.20) allows us to definitively connect the differential operator $\mathcal{A}$ (defined precisely in Proposition 5) governing (3.37) to the dispersal dynamics of overcrowding for two species. As to the connection with systems of type (1.2) and (2.18) commonly use din the literature, we de believe that is likely more nuanced and may hold true in specific cases, particularly when solutions exist, such as in fully segregated scenarios.
@@
    Theorem 2 also yields other crucial results which uses the equivalence between weak solutions and $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow solutions. We focus here on the uniqueness and large time behavior which could be of great importance for the applications :
    1. (a)
      Uniqueness : Since, in the case where $V\equiv 0,$ a weak solution to equation (3.37) corresponds to the strong solution of (2.21), we deduce that they are unique. Notably, this allows to infer the uniqueness of weak solutions (whenever they exist) in fully segregated scenarios.
    2. (b)
      Asymptotic behavior $t\to\infty$ : thanks to the theory of gradient flows in Hilbert spaces, we see that if $f\in L^{2}(0,\infty;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)),$ satisfies
      $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-\lim_{t\to\infty}f(t)=f_{\infty},$
      then the solution $\rho(t)$ converges uniformly in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)$ to $\rho,$ as $t\to\infty,$ where $\rho$ is such that there exists $\eta\in[H^{1}(\Omega)]^{2}$ for which the couple $(\rho,\eta)$ is a solution of the stationary problem
      (2.22) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})=f_{\infty 0k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})\cdot\nu=\pi_{k}&\hbox{ on}\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}}.\end{array}\right.$
@@
    In the case where $V\not\equiv 0,$ the $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow theory is no longer applicable. The uniqueness for weak solutions of the problem (3.37) remains to be a challenging open problem. Remember for the case of one specie the uniqueness may follows through contraction principles in $L^{1}$ and/or in $\mathcal{W}_{2}-$Wasserstein space (one can see [18, 22] for discussions and references in this direction). It is not clear yet for us how to deal with $L^{1}-$theory for cross-diffusion of the type (3.37).
@@
    We focus here in the case of source term $f=f(t,x).$ However, one sees that one can use the results of Theorem 2 also for treating more general reaction terms $f=f(t,x,\rho,\eta).$

### v2

@@ 2.1 Existence result
- Parabolic nature: In the aggregated region, $S,$ the systems diverge significantly. While (2.15) exhibits a fully parabolic regime with respect to all three variables $\rho_{1},$ $\rho_{2}$ and $\rho_{1}+\rho_{2},$ equation (2.21) is only parabolic with respect $\rho_{1}+\rho_{2}.$
- Correlation of solutions in the segregation regime: We observe that if a solution $(\rho,\eta)$ to (2.15) satisfies the condition $[\rho_{1}\rho_{2}\neq 0]=\emptyset,$ (i.e., purely segregation regime), then it is also a weak solution to (2.21). However, it is not yet clear if the converse implication holds true in general. The existence of a segregation regime for the system described by Equation (2.15) is an intriguing question that we do not address in this paper.
@@
Since we’re only interested in nonnegative solutions, the behavior of the function $\beta(x,r)$ for negative values of $r$ is irrelevant. We can assume that $\beta(x,r)=0$ for any $r\leq 0.$ As a consequence, for any $s\in\partial\beta(x,r),$ we have $s\geq 0.$ This nonnegativity property has a crucial implication for solutions of equation (2.15). Specifically, for any solution $(\rho,\eta)$ we can systematically conclude that:
(2.22) $\eta_{1}\vee\eta_{2}\geq 0\quad\text{ a.e. in }Q.$
2.2 $H^{-1}-$gradient flow approach
@@
(2.23) $\|f_{k}\|_{H^{-1}_{\Gamma_{D_{k}}}}=\min_{\omega\in L^{2}(\Omega)^{N}}\left\{\left(\int_{\Omega}\sigma_{k}\>|\omega|^{2}\right)^{1/2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
@@
$\langle f_{k},g_{k}\rangle_{k,\Omega}:=\int_{\Omega}\sigma_{k}\>\phi^{f}_{k}\cdot\phi^{g}_{k},\quad\hbox{ for any }f,g\>\in H^{-1}_{\Gamma_{D_{k}}}(\Omega)$
where $\phi^{f}_{k}$ and $\phi^{g}_{k}$ are (unique) given by
$\phi^{f}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}\sigma_{k}\>|\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega+\overline{f}_{k})=f_{0k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega+\overline{f}_{k})\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}$
@@
This being said, let us consider the space $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),$ endowed with the norm
@@
This is in turn a Hilbert space, and the associate inner product is given by
@@
(2.24) $\mathcal{E}(\rho):=\left\{\begin{array}[]{ll}\int_{\Omega}\beta(.,\rho_{1}+\rho_{2})&\hbox{ if }\rho_{1},\>\rho_{2},\>\beta(.,\rho_{1}+\rho_{2})\in L^{1}_{+}(\Omega)\\ \\ +\infty&\hbox{ otherwise }.\end{array}\right.$
Thanks to assumptions $(H1)$ and $(H2)$, it is not difficult to see that $\mathcal{E}$ is convex and l.s.c. in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega).$ This implies that the sub-differential of $,\displaystyle\mathcal{E},$$\partial\mathcal{E},$ is well-defined and forms a maximal monotone graph within $\left(H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega),\|-\|_{H^{-1}_{\Gamma_{D}}}\right).$ So, we can consider the $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow problem
(2.25) $\left\{\begin{array}[]{ll}\frac{d}{dt}\rho(t)+\partial\mathcal{E}(\rho(t))\ni h(t)&\hbox{ a.e. }t\in(0,T)\\ \\ \rho(0)=\rho_{0},\end{array}\right.$
where $h\in L^{2}(0,T;H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega)).$ Our aim now, is to establish the connection between the problems (2.15) and (2.25). This is summarized in the following theorem.
@@
Under the assumptions $g\equiv 0$, $\pi\equiv 0$, $f\in L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\right)$ and $V\in[L^{\infty}(Q)]^{2},$ let us consider $(\rho,\eta)$ be a weak solution of the system of PDE (2.15) given by Theorem 1. Then, $\rho$ coincides with the strong solution of (2.25), where
$h=(f_{1}-\nabla\cdot(\rho_{1}\>V_{1}),f_{2}-\nabla\cdot(\rho_{2}\>V_{2})).$
Thanks to general theory of evolution problem governed by maximal monotone graph on Hilbert space, this theorem implies in particular that, if $(\rho^{1},\eta^{1})$ and $(\rho^{2},\eta^{2})$ are weak solutions of the system of PDE (2.15) corresponding to given $(f^{1},V^{1})$ and $(f^{2},V^{2}),$ respectively, then
(2.26) $\begin{array}[]{c}\frac{1}{2}\frac{d}{dt}\sum_{k=1,2}\|\rho^{1}_{k}-\rho^{2}_{k}\|_{H^{-1}_{\Gamma_{D_{k}}}(\Omega)}^{2}+[\partial\mathcal{E}(\rho^{1})-\partial\mathcal{E}(\rho^{2}),\rho^{1}-\rho^{2}]_{\Omega}\\ \leq\sum_{k=1,2}\langle f^{1}_{k}-f^{2}_{k},\rho^{1}_{k}-\rho^{2}_{k}\rangle_{k,\Omega}-\sum_{k=1,2}\langle\nabla\cdot(\rho^{1}_{k}\>V^{1}_{k}-\rho^{2}\>V^{2}_{k}),\rho^{1}_{k}-\rho^{2}_{k}\rangle_{k,\Omega},\quad\hbox{ in }\mathcal{D}^{\prime}(0,T).\end{array}$
More specifically, this implies the following uniqueness result.
Corollary 1
Under the assumptions $g\equiv 0$ $\pi\equiv 0$ and $f\in L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\right)$, let us consider $V^{1},\>V^{2}\in L^{\infty}(Q)$. If $(\rho^{1},\eta^{1})$ and $(\rho^{2},\eta^{2})$ are weak solutions of the system of PDE (2.15) corresponding to $V^{1}$ and $V^{2},$ respectively, such that $\rho^{1}(0)=\rho^{2}(0)$ and there exists $c>0$ such that
(2.27) $\begin{array}[]{c}[\partial\mathcal{E}(\rho^{1})-\partial\mathcal{E}(\rho^{2}),\rho^{1}-\rho^{2}]_{\Omega}+\sum_{k=1,2}\langle\nabla\cdot(\rho^{1}_{k}\>V^{1}_{k}-\rho^{2}\>V^{2}_{k}),\rho^{1}_{k}-\rho^{2}_{k}\rangle_{k,\Omega}\\ \geq-c\>\sum_{k=1,2}\|\rho^{1}_{k}-\rho^{2}_{k}\|_{H^{-1}_{\Gamma_{D_{k}}}(\Omega)}^{2},\quad\hbox{ a.e. in }(0,T),\end{array}$
then $\rho_{1}=\rho_{2}.$
Proof :
The proof is a simple consequence of (2.26) and Grönwall inequality. $\square$
Equation (2.27) incorporates assumptions about $\beta$ and $V^{s},$ for $s=1,2.$ This condition becomes particularly relevant in applications where the drift term depends on the densities themselves. However, in the special case where $V\equiv 0$ (i.e., no potential term), the uniqueness of the weak solution can be established without additional assumptions. This directly follows from the monotone property of $\partial\mathcal{E}$, expressed as :
$[\partial\mathcal{E}(\rho^{1}(t))-\partial\mathcal{E}(\rho^{2}(t)),\rho^{1}-\rho^{2}]_{\Omega}\geq 0.$
Therefore, the following uniqueness result holds.
Corollary 2
Under the assumptions $g\equiv 0,$ $\pi\equiv 0$, $V\equiv 0,$ $f\in L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)\right)$ and $\rho_{0}\in L^{2}(Q)\times L^{2}(Q)$ satisfying (2.16), the system of PDE (2.15) admits a unique weak solution.
@@
    Theorem 2 bridges a connection between the cross-diffusion system (3.51) and the well-established theory of gradient flows in Hilbert spaces. Furthermore, the specific form of the internal energy (2.24) allows us to definitely connect the differential operator $\mathcal{A}$ (defined precisely in Proposition 5) governing (3.51) to the dispersal dynamics of overcrowding for two species. As to the connection with systems of type (1.2) and (2.21) commonly used in the literature, we de believe that is likely more nuanced and may hold true only in specific cases, particularly when solutions exist, such as in fully segregated scenarios.
@@
    Theorem 2 also yields other crucial results which uses the equivalence between weak solutions and $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow solutions. Beyond uniqueness results of Corollary 1 and Corollary 2, the large time behavior which could be of great importance for the applications maybe concerned too. Indeed, thanks to the theory of gradient flows in Hilbert spaces, we see that if $f\in L^{2}(0,\infty;H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega))$ is such that, as $t\to\infty,$
    $f(t)\to f_{\infty},\quad\hbox{ in }H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}},$
    then the solution $\rho(t)$ converges uniformly in $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)$ to $\rho,$ as $t\to\infty,$ where $\rho\in L^{2}(\Omega)\times L^{2}(\Omega)$ satisfies (2.16) and there exists $\eta\in[H^{1}(\Omega)]^{2}$ for which the couple $(\rho,\eta)$ is a solution of the stationary problem
    (2.28) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})=f_{\infty 0k},\quad\hbox{ for }k=1,2\\ \\ \eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2}),\>\quad\eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{\infty k})\cdot\nu=\pi_{k}&\hbox{ on}\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}}.\end{array}\right.$
@@
    In the case where $V\not\equiv 0,$ the application of $H^{-1}_{\Gamma_{D_{1}}}\!\!(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}\!\!(\Omega)-$gradient flow theory remains to be restrictive. The uniqueness of weak solutions of the problem (3.51) remains to be a challenging open problem. Remember that for the case of one specie, the uniqueness may follows through contraction principles in $L^{1}$ and/or in $\mathcal{W}_{2}-$Wasserstein space (one can see [17, 21] for discussions and references in this direction). It is not clear yet for us how to deal with $L^{1}-$theory for cross-diffusion of the type (3.51).
@@
    In this paper, we focus on the case where source term $f=f(t,x).$ However, one sees that one can use the results of Theorem 2 also for treating more general reaction terms $f=f(t,x,\rho,\eta).$
@@
2.3 Weighted internal energy
The results of Section 2.1 and Section 2.2 enable also to treat the case where the densities are combined with specific weights in the internal energy. These weights represent the relative importance of each density in influencing the internal energy. In this case the cross-diffusion system reads
(2.29) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})=f_{0k},\\ \\ \eta_{k}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})+\partial I\!\!I_{[0,\infty)}(\rho_{k})\end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\rho_{k}V_{k}+\overline{f}_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\\ \\ \rho_{k}(0)=\rho_{0k}&\hbox{ in }\Omega,\end{array}\right\}\hbox{ for }k=1,2,$
where $(\alpha_{1},\alpha_{2})\in I\!\!R^{2}$ are given and assumed to satisfy the following condition
(2.30) $0<\min(\alpha_{1},\alpha_{2}).$
Theorem 3
Under the previous assumptions, assume moreover that
$\int_{\Omega}\beta(.,\alpha_{1}\rho_{01}+\alpha_{2}\rho_{02})<\infty.$
Then, the system of PDE (2.29) has a solution $(\rho,\eta)$ in the sense that, for each $k=1,2,$ $\rho_{k}\in L^{2}_{+}(Q)\cap W^{1,2}(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)),$ $\rho_{k}(0)=\rho_{0k},$ $\eta_{k}\in L^{2}(0,T;H^{1}(\Omega)),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
(2.31) $\left\{\begin{array}[]{l}\eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),\\ \\ \eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\end{array}\right.\quad\hbox{ a.e. in }Q,$
and
(2.32) $\frac{d}{dt}\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\rho\>V_{k}+\overline{f}_{k})\cdot\nabla\xi_{k}=\int_{\Omega}f_{0k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{N_{k}}},\quad\hbox{ in }\mathcal{D}(^{\prime}0,T),$
Proof :
The proof of this result follows directly by using Theorem 1 and the fact that $(\rho,\eta)$ is a solution of (2.29) if and only if the couple $(\tilde{\rho},\eta),$ where $\tilde{\rho}:=(\alpha_{1}\>\rho_{1},\alpha_{2}\>\rho_{2}),$ is a solution of the
(2.33) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\tilde{\rho}_{k}-\nabla\cdot(\sigma_{k}/\alpha_{k}\>\nabla\eta_{k}+\tilde{\rho}_{k}V_{k}+\overline{f}_{k}/\alpha_{k})=f_{0k}/\alpha_{k},\\ \\ \eta_{k}\in\partial\beta(x,\tilde{\rho}_{1}+\tilde{\rho}_{2})+\partial I\!\!I_{[0,\infty)}(\tilde{\rho}_{k})\end{array}\right\}&\hbox{ in }Q,\\ \\ (\sigma_{k}/\alpha_{k}\>\nabla\eta_{k}+\tilde{\rho}_{k}V_{k}+\overline{f}_{k}/\alpha_{k})\cdot\nu=\pi_{k}/\alpha_{k}&\hbox{ on }\Sigma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Sigma_{D_{k}},\\ \\ \rho_{k}(0)=\rho_{0k}&\hbox{ in }\Omega,\end{array}\right\}\hbox{ for }k=1,2.$
$\square$